(a) Evaluate∫[


𝒙/(𝒙^2+𝟏)^(1/2)𝒅𝒙.


(b) Use MATLAB to generate some typical integral curves of 𝑓(π‘₯) =


𝒙/(𝒙^2+𝟏)^(1/2)𝒅𝒙over the interval (βˆ’5,5).



1
Expert's answer
2022-01-06T11:44:37-0500

Solution:

(a):

"I=\int \dfrac{x}{\sqrt{x^2+1}}dx"

"I=\int \dfrac{x}{\sqrt{x^2+1}}dx"

Put "x^2+1=t"

"\Rightarrow 2xdx=dt
\\\Rightarrow xdx=\dfrac{dt}{2}"

So, "I=\dfrac{1}{2}\int \dfrac{1}{\sqrt{t}}dt"

"I=\dfrac{1}{2}\int \dfrac{1}{\sqrt{t}}dt"

"I=\dfrac{1}{2}\int t^{-1/2}dt
\\=\dfrac{1}{2}.\dfrac{t^{1/2}}{1/2}+c
\\=\sqrt t+c
\\=\sqrt{x^2+1}+c"

"I=\dfrac{1}{2}\int t^{-1/2}dt
\\=\dfrac{1}{2}.\dfrac{t^{1/2}}{1/2}+c
\\=\sqrt t+c
\\=\sqrt{x^2+1}+c""I=\dfrac{1}{2}\int t^{-1/2}dt
\\=\dfrac{1}{2}.\dfrac{t^{1/2}}{1/2}+c
\\=\sqrt t+c
\\=\sqrt{x^2+1}+c"

(b):

The graph using matlab from (-5,5) is as follows: